OpenStudy (jazzyfa30):

Factor Completely: a2+ 5a +ab Select one: a. (a+5)(a+b) b. a(a+5+b) c. (a+2)(a+3) d. prime

OpenStudy (jazzyfa30):

@mathstudent55

OpenStudy (anonymous):

C

thomaster (thomaster):

We have \(\large\sf a^2+5a+ab\) As you can see, a is in all the 3 parts so we can put that one outside of the parenthesis. \(\sf a*a=a^2\) \(\sf a*5=5a\) \(\sf a*b=ab\) \(\large\sf a(a+5+b)=a^2+5a+ab\)

OpenStudy (anonymous):

i mean B